# FML-bench: A Benchmark for Automatic ML Research Agents Highlighting the Importance of Exploration Breadth

### v1

@@ Abstract
Large language models (LLMs) have sparked growing interest in automatic machine learning research agents. Among them, agents capable of autonomously proposing ideas and conducting machine learning experiments are particularly promising, as they maximize research automation and accelerate scientific progress by iteratively refining ideas based on experimental results. However, comprehensively evaluating such agents remains challenging. Existing benchmarks tend to overemphasize engineering aspects while neglecting academic rigor, creating barriers that obscure a clear assessment of an agent’s scientific capabilities in machine learning research. They also suffer from limited task diversity, an overemphasis on application-oriented tasks over fundamental research problems, and limited scalability to realistic research settings. To address these limitations, we introduce FML-bench, a benchmark designed to evaluate automatic machine learning research agents on 8 diverse and fundamental machine learning research problems. It reduces coding burden, emphasizes fundamental problems rather than specific use cases, offers high task diversity, and is extensible to real-world machine learning GitHub repositories. Furthermore, we present a unified evaluation framework with five complementary metrics, designed to comprehensively assess agent performance on our benchmark. We evaluate state-of-the-art automatic research agents on FML-bench, and find that agents employing broad research exploration strategies outperform those focusing on narrow but deep exploration. These findings suggest that emphasizing the breadth of exploration may lead to more effective research outcomes than focusing solely on incremental refinement. Our benchmark is available at: https://github.com/qrzou/FML-bench.
footnotetext: * Equal contribution.  $\dagger$ Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have catalyzed a resurgence of interest in automatic machine learning (ML) research agents that assist or carry out parts of the scientific discovery workflow. These agents not only support hypothesis generation, coding, and experiment management, but also increasingly act as collaborators in discovery by providing complementary perspectives that can accelerate machine learning research across domains. Within this landscape, agents that automatically propose ideas and run experiments are particularly compelling (Lu et al., 2024; Yamada et al., 2025). They close the loop from ideation to empirical validation to maximize automation and to speed up research cycles. Compared to settings that only elicit ideas and then use LLMs or humans to assess “novelty” and “feasibility” which often diverge from real-world utility, this approach evaluates agents based on actual experimental outcomes, providing objective and quantitative evidence of their effectiveness (Wang et al., 2024a; Baek et al., 2024; Si et al., 2024).
Despite rapid progress, existing benchmarks offer an incomplete picture of research competence, as shown in Tab. 1. Most focus on Kaggle-style, use-case-oriented tasks and emphasize engineering execution (e.g., feature engineering, standardized model training, and optimization) while paying limited attention to evaluating an agent’s ability to tackle fundamental machine learning research problems, such as representation learning and generalization (Chan et al., 2024; Huang et al., 2023; Padigela et al., 2025; Jing et al., 2024). Moreover, some benchmarks provide only raw data without baseline code (Chan et al., 2024; Jing et al., 2024), making it difficult to systematically assess agents’ research capabilities while introducing coding barriers that can obscure academic merit (e.g., when sound ideas fail due to engineering pitfalls). In addition, even when baseline codebases are provided, they are often handcrafted and tightly formatted (Huang et al., 2023; Padigela et al., 2025), which hinders their scalability. Because adapting them to new tasks usually requires substantial re-engineering to fit their benchmark design, rather than allowing direct use of existing codebases.
@@
To address these gaps, we introduce FML-bench, a benchmark designed to evaluate automatic ML research agents on fundamental ML problems. FML-bench comprises 8 diverse tasks (Fig. 1) chosen to reflect bottlenecks that repeatedly surface in modern ML. The design follows four principles. 1) Fundamental ML problems. Designed tasks target core scientific challenges rather than application products or leaderboard scoring, keeping the focus on research questions. 2) Real-world codebases. Tasks are instantiated from existing research repositories, mirroring typical practice where new ideas are tested by adapting prior code. 3) Extensibility by construction. The benchmark can easily incorporate machine learning GitHub repositories that support end-to-end training and evaluation, requiring only minor output-format adapters. 4) Low coding barrier. Agents are not required to build entire codebases from scratch, but can start from provided baselines. This setup enables agents to focus on scientific advances in algorithms and architectures rather than on purely engineering effort.
The tasks included in FML-bench span a broad set of foundational problems (Wang et al., 2022; Pearl, 2019; Adadi, 2021; Goodfellow et al., 2014; Abadi et al., 2016; Bengio et al., 2013; Chen & Liu, 2018; Mehrabi et al., 2021): generalization (cross-domain transfer), data efficiency (learning from few samples), representation learning (discovery of meaningful features), continual learning (knowledge retention over time), causality (intervention reasoning), robustness and reliability (resistance to adversarial corruption), privacy (protection against information leakage) and fairness and bias (equitable treatment across groups). Agents are expected to propose new or improved ML methods that deliver stronger empirical results than baselines across these eight tasks.
To evaluate agents holistically, we formalize five complementary metrics that capture different facets of research competence. Utility measures empirical performance improvement and serves as the primary objective. Diversity quantifies the variety of code modifications proposed; empirically, broader exploration often correlates with larger gains. Academic contribution rate distinguishes academic modifications (e.g., new losses, architectures, or training schemes) from engineering modifications (e.g., hyperparameter tuning), rewarding agents that prioritize scientifically meaningful changes. Cost accounts for computational and time expenditure. Step success rate captures the fraction of runs that produce valid results without bugs, reflecting an agent’s reliability in multi-step workflows. Different tasks emphasize different subsets of these metrics so that the aggregate evaluation reflects the spectrum of capabilities required for fundamental ML research.
We evaluate several state-of-the-art automatic research agents and LLMs on FML-bench. A central finding concerns agent strategy: once the basic requirements for both exploration breadth and depth are met, broader exploration proves more effective. Generating a wider variety of ideas more reliably leads to successful methods than repeatedly refining a single one, and we observe a positive correlation between idea diversity and performance improvement. Besides, we find that Gemini-2.5-Pro outperforms GPT-5 under our protocol. Finally, while CLI-style agents such as Claude Code offer general-purpose flexibility, they often fail to complete multi-step tasks due to early termination, where the model stops despite further actions being possible. This suggests that, although flexible, CLI-style agents are less suitable for automatic machine learning research than agents specifically designed for it.
We summarize our contributions as follows:
  • We construct FML-bench, a benchmark centered on a diversity of fundamental ML problems instantiated in real-world codebases, closing gaps left by use-case-oriented, engineering-heavy evaluations.
  • We propose a five-dimensional evaluation protocol covering utility, diversity, academic contribution rate, cost, and step success rate, jointly measuring empirical progress, research quality, and reliability.
  • We provide empirical insights on research strategy (breadth vs. depth of exploration), quantify the role of diversity in driving gains, and report comparative results across leading agent frameworks and LLMs, offering guidance for practical agent design.
@@
Table 1: Comparison of ML agent benchmarks across key design goals and agent requirements. Repo refers to the repository, and Comp denotes Competition. : In MLAgentBench, only part of the tasks meet this requirement; users must prepare baseline and evaluation code even when some tasks are based on real-world Kaggle repositories.
Criterion Ours MLE–Bench MLAgentBench ML–Dev–Bench DSBench
Design Goals
@@
Scalability via Existing Repo/Comp
Requirements to Agent
Understand Codebase
Understand Data
Execute Arbitrary CMD with Args
Execute Multi-step CMD List
@@ 2 Related Works
2.1 Automatic AI Agents
Recent advances in large language models (LLMs) have enabled research agents to support core components of the scientific workflow. These agents are capable of generating and prioritizing research ideas, retrieving and synthesizing literature, and simulating peer review processes. For instance, SciMON (Wang et al., 2024a) and Nova (Hu et al., 2024) implemented frameworks for generating diverse and novel research ideas. AutoSurvey (Wang et al., 2024b) presented an automated literature review framework that performs retrieval over a large arXiv corpus, followed by outline planning and section drafting using specialized models. Meanwhile, AgentReview (Jin et al., 2024) employed LLM agents to simulate peer reviews, rebuttals, and committee discussions, offering insights into the dynamics of academic decision-making.
@@ 2.2 Benchmarks for ML Agents
While recent benchmarks have begun to evaluate agents on code-intensive tasks, they remain limited in both scope and flexibility. MLAgentBench (Huang et al., 2023) includes 13 machine learning engineering tasks, but most are implemented as single-file scripts, which is not practical for real-world scenarios. In addition, it requires setting an individual evaluator for each task and lacks support, limiting its scalability to support more tasks. MLE-Bench (Chan et al., 2024) covers 75 Kaggle competitions and assesses whether agents can function as machine learning engineers. It emphasizes tasks such as data pipeline management, experiment orchestration, and submission formatting, which may shift focus away from core machine learning understanding. ML-Dev-Bench (Padigela et al., 2025) places greater emphasis on engineering aspects such as dataset loading and API integration. It evaluates agents’ ability to improve existing baselines only in performance tests, which are relatively simple due to narrow task scopes like classification and segmentation, and the use of fixed starter files. In contrast, our benchmark includes tasks spanning diverse machine learning domains. DSBench (Jing et al., 2024) aggregates 466 data analysis tasks and 74 modeling tasks from ModelOff and Kaggle, focusing on problem-solving within data science workflows. By comparison, our benchmark focuses on 8 diverse and fundamental machine learning research problems. It is built on real-world codebases, thereby providing practical challenges and strong extensibility by construction, while maintaining a low coding barrier.
3 Unified Evaluation Framework for Automatic ML Research
Automatic ML research agents operate through iterative refinement cycles, where each iteration involves hypothesis generation, code modification, experimental execution, and empirical evaluation. To formalize this process, we propose a unified optimization framework that is explicitly aligned with five evaluation dimensions.
@@ 3 Unified Evaluation Framework for Automatic ML Research
Consider an agent conducting research over $T$ iterations. At iteration $t\in\{1,\ldots,T\}$, the agent starts with codebase $\mathbf{C_{t-1}}$ and generates a hypothesis $h_{t}$ from a learned proposal distribution $q_{t}$. This hypothesis is instantiated as a concrete code modification $m_{t}$. After completing all $T$ iterations, the agent has produced a hypothesis set $\mathcal{H}=\{h_{1},\ldots,h_{T}\}$ with corresponding modifications $\mathcal{M}=\{m_{1},\ldots,m_{T}\}$. The agent’s objective over the research process is:
$\max_{T,\,\{q_{t},m_{t}\}_{t=1}^{T}}\ \sum_{t=1}^{T}\left[U_{t}+\lambda A_{t}-\eta P_{t}\right]+\gamma S(\mathcal{M},\mathbf{C_{1}})+\beta D(\mathcal{H})$ (1)
where $U_{t}=\mathbb{E}_{h_{t}\sim q_{t}}[U(m_{t},\mathbf{C_{t-1}})]$ is the expected utility, $A_{t}=A(m_{t})$ is the academic contribution rate, $\mathbf{C_{t}}=\mathbf{C_{t-1}}\oplus m_{t}$, $P_{t}=P(m_{t})$ is the cost, $S(\mathcal{M},\mathbf{C_{1}})$ is the step success rate, and $D(\mathcal{H})$ measures diversity across all hypotheses.
Evaluation metrics
Each term in Eq. 1 corresponds directly to our proposed evaluation metrics:
  • Utility $U(m,\mathbf{C})$: The primary objective measuring empirical performance improvement. Specifically, $U(m,\mathbf{C})=\text{perf}(\mathbf{C}\oplus m)-\text{perf}(\mathbf{C})$, where $\text{perf}(\cdot)$ evaluates the task-specific metric (e.g., accuracy, AUC, error rate).
  • Diversity $D(\mathcal{H})$: Quantifies the variety across all proposed hypotheses $\mathcal{H}=\{h_{1},\ldots,h_{T}\}$. We measure this through semantic and structural variance of the resulting modifications, capturing the agent’s exploration breadth. Empirically, $D(\mathcal{H})$ strongly correlates with discovering high-performing solutions.
  • Academic Contribution Rate $A(m)$: Measures the proportion of academic/algorithmic contributions (e.g., novel architectures, loss functions, training schemes) relative to engineering modifications (e.g., hyperparameter tuning, infrastructure fixes). Higher $A(m)$ indicates greater scientific contribution, distinguishing genuine research advances from implementation optimizations.
@@
  • Step Success Rate $S(\mathcal{M},\mathbf{C_{1}})$: Captures the reliability of all code modifications $\mathcal{M}$ on initial codebase $\mathbf{C_{1}}$. This reflects the agent’s ability to produce syntactically correct, semantically coherent code that successfully completes experiment iterations without errors.
  • Cost $P(m)$: Encompasses time expenditure (wall-clock time) and API usage (tokens) required to execute codebase with modification $m$.
Design Principles for Effective Agents.
To achieve high utility while maintaining research quality, effective agents should satisfy:
$D(\mathcal{H})\geq\delta,\quad\bar{A}\geq\alpha,\quad S\geq\rho,\quad\sum_{t=1}^{T}P_{t}\leq B$ (2)
where $\bar{A}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$ is the average academic contribution rate. These principles guide agent design: maintain exploration breadth to avoid local optima ($D\geq\delta$), prioritize algorithmic innovation ($\bar{A}\geq\alpha$), ensure reliable execution ($S\geq\rho$), and respect computational budgets ($\sum P_{t}\leq B$).
4 Benchmark Design
4.1 Task Description
We select 8 diverse tasks that collectively span the most critical aspects of ML research, ensuring a comprehensive assessment for ML research agents. These tasks represent core competencies that a comprehensive and robust agent should demonstrate: the ability to generalize beyond training distributions, learn efficiently from limited data, discover meaningful representations, retain knowledge over time, reason about causal relationships, maintain reliability under adversarial conditions, protect sensitive information, and operate fairly across different groups.
Specifically, we evaluate agent performance across 8 critical machine learning tasks. 1) Generalization. Assessed via a cross-domain transfer task in which models train on a source domain and are evaluated on a held-out target under distribution shift. The objective is to maximize out-of-domain accuracy. 2) Data Efficiency. Tested through a few-shot classification task. Agents should propose approaches to improve metric-based decision rules in the embedding space to boost accuracy with limited labels. 3) Representation Learning. Pretrain encoders in a self-supervision manner and evaluate by linear-probe accuracy with the encoder frozen, targeting meaningful feature discovery. 4) Continual Learning. Measure knowledge retention in a class-incremental sequence with a shared output head. Agents should propose methods to mitigate catastrophic forgetting and maximize average accuracy across all tasks. 5) Causality. Estimate treatment effects under a specified causal data-generating process and minimize absolute error in the average treatment effect (ATE). 6) Robustness and Reliability. Evaluate resilience to adversarial corruption, including poisoning or backdoor perturbations, while preserving clean performance. The defense score balances both objectives. 7) Privacy. Assess protection against information leakage by reducing the effectiveness of membership-inference attacks, i.e., lowering attack AUC. 8) Fairness and Bias. Evaluate equitable performance across groups in binary classification with sensitive attributes, aiming to improve group-fairness metrics (e.g., minimizing absolute average odds difference) without sacrificing overall accuracy.

### v2

# FML-bench: Benchmarking Machine Learning Agents for Scientific Research

@@ Abstract
Large language models (LLMs) have sparked growing interest in machine learning research agents that can autonomously propose ideas and conduct experiments. However, existing benchmarks predominantly adopt an engineering-oriented perspective: they emphasize application-oriented tasks and evaluate primarily on final performance and computational cost, overlooking agents’ research processes and limiting assessment of their capabilities in scientific research settings. To more comprehensively evaluate agents in scientific research settings, we introduce FML-bench, a benchmark comprising 8 diverse and fundamental ML research tasks, and further propose complementary metrics, notably Exploration Diversity, which quantifies the variance of proposals across iterations and reveals how exploration patterns influence research outcomes. We evaluate state-of-the-art research agents on FML-bench, showing that agents employing broad exploration strategies exhibit higher exploration diversity and achieve superior performance, and that exploration diversity positively correlates with performance improvements across multiple tasks. We hope these findings and our benchmark inform future agent design and support the community in further investigating agent behavior. Our benchmark is available at: https://github.com/qrzou/FML-bench.
Machine Learning, Large Language Models, Research Agents, Benchmarking, Evaluation, Autonomous Research, Exploration Diversity, ICML
@@ 1 Introduction
Large language models (LLMs) have catalyzed a resurgence of interest in machine learning (ML) research agents which assist or carry out parts of the scientific discovery workflow. These agents not only support hypothesis generation, coding, and experiment management, but also increasingly act as collaborators in discovery by providing complementary perspectives that can accelerate machine learning research across domains. Within this landscape, agents that automatically propose ideas and run experiments are particularly compelling (Lu et al., 2024; Yamada et al., 2025). They close the loop from ideation to empirical validation to maximize automation and to speed up research cycles. Compared to settings that only elicit ideas and then use LLMs or humans to assess “novelty” and “feasibility” which often diverge from real-world utility, this approach evaluates agents based on actual experimental outcomes, providing objective and quantitative evidence of their effectiveness (Wang et al., 2024a; Baek et al., 2024; Si et al., 2024).
Despite rapid progress, existing benchmarks offer an incomplete picture of research competence, as shown in Tab. 1. These benchmarks predominantly adopt an engineering-oriented perspective rather than one aligned with scientific research settings, and this limitation manifests in two key aspects. First, regarding task construction, most benchmarks focus on Kaggle-style, application-oriented tasks that emphasize engineering execution (e.g., feature engineering, standardized model training, and optimization) while paying limited attention to evaluating an agent’s ability to tackle fundamental machine learning research problems, such as representation learning and generalization (Chan et al., 2024; Huang et al., 2023; Padigela et al., 2025; Jing et al., 2024). Second, regarding evaluation design, these benchmarks primarily assess final task performance metrics (e.g., accuracy, recall) and computational cost, while overlooking the characteristics of agents’ internal iterative processes, which limits analysis of how agent characteristics relate to research outcomes.
@@
To address these gaps and more comprehensively evaluate agents in scientific research settings, we introduce FML-bench, a benchmark designed to assess automatic ML research agents on fundamental ML problems. FML-bench comprises 8 diverse tasks (Fig. 1) chosen to reflect bottlenecks that repeatedly surface in modern ML. The tasks span eight diverse and fundamental ML research problems, including generalization, data efficiency, representation learning, continual learning, causality, robustness, privacy, and fairness (see Section 3 for task design rationale and references). Agents are expected to propose new or improved ML methods that deliver stronger empirical results than baselines across these tasks.
Moreover, to enable deeper analysis of agent behavior beyond final outcomes, we further propose complementary metrics that characterize the research process. Notably, we introduce Exploration Diversity, which quantifies the variance of proposals across iterations and reveals how exploration patterns influence research outcomes. Specifically, we compute code embeddings using GraphCodeBERT (Guo et al., 2020) and measure the dispersion of these embeddings around their centroid. Additionally, we propose Step Success Rate and Step Completion Rate to characterize agent reliability throughout the research process. We also report commonly used metrics, including Performance and Cost.
Beyond the two core contributions above, FML-bench addresses additional limitations of existing benchmarks. Some benchmarks provide only raw data without baseline code (Chan et al., 2024; Jing et al., 2024), making it difficult to systematically assess agents’ research capabilities while introducing coding barriers that can obscure academic merit (e.g., when sound ideas fail due to engineering pitfalls). Even when baseline codebases are provided, they are often handcrafted and tightly formatted (Huang et al., 2023; Padigela et al., 2025). In contrast, FML-bench tasks are constructed upon widely recognized ML research repositories with carefully designed task specifications, reflecting how researchers typically build upon established codebases to investigate new ideas. Furthermore, agents are not required to build entire codebases from scratch but can start from provided baselines, enabling them to focus on scientific advances in algorithms and architectures rather than purely engineering effort.
@@
Table 1: Comparison of ML agent benchmarks across key design goals. Repo refers to the repository, and Comp denotes Competition. : In MLAgentBench, only part of the tasks meet this requirement; users must prepare baseline and evaluation code even when some tasks are based on real-world Kaggle repositories.
Design Goals Ours MLE–Bench MLAgentBench ML–Dev–Bench DSBench
@@
We evaluate several state-of-the-art automatic research agents on FML-bench. The results show that agents employing broad exploration strategies exhibit higher exploration diversity and achieve superior performance, and that exploration diversity positively correlates with performance improvements across multiple tasks.
We summarize our contributions as follows:
  • We construct FML-bench, a benchmark for scientific research centered on diverse fundamental ML problems, closing gaps left by application-oriented, engineering-heavy evaluations.
  • We propose complementary metrics for analyzing agent behavior throughout the scientific research process, notably Exploration Diversity, which quantifies the variance of proposals across iterations, along with Step Success Rate and Step Completion Rate for assessing agent reliability.
  • We report empirical findings suggesting that broad exploration strategies and high exploration diversity are important contributors to effective ML research agents.
@@ 2 Related Works
2.1 Automatic Research Agents
With the emergence of large language models (LLMs), research agents have increasingly been explored as tools to support core components of the scientific workflow. These agents are capable of generating and prioritizing research ideas, retrieving and synthesizing literature, and simulating peer review processes. For instance, SciMON (Wang et al., 2024a) and Nova (Hu et al., 2024) implemented frameworks for generating diverse and novel research ideas. AutoSurvey (Wang et al., 2024b) presented an automated literature review framework that performs retrieval over a large arXiv corpus, followed by outline planning and section drafting using specialized models. Meanwhile, AgentReview (Jin et al., 2024) employed LLM agents to simulate peer reviews, rebuttals, and committee discussions, offering insights into the dynamics of academic decision-making.
@@ 2.2 Benchmarks for ML Agents
Existing benchmarks have begun to evaluate agents on code-intensive tasks, yet they remain limited in both scope and flexibility. MLAgentBench (Huang et al., 2023) includes 13 machine learning engineering tasks, but most are implemented as single-file scripts, which is not practical for real-world scenarios. In addition, it requires to set individual evaluator for each task and lacks support, limiting its scalability to support more tasks. MLE-Bench (Chan et al., 2024) covers 75 Kaggle competitions and assesses whether agents can function as machine learning engineers. It emphasizes tasks such as data pipeline management, experiment orchestration, and submission formatting, which may shift focus away from core machine learning understanding. ML-Dev-Bench (Padigela et al., 2025) places greater emphasis on engineering aspects such as dataset loading and API integration. It evaluates agents’ ability to improve existing baselines only in performance tests, which are relatively simple due to narrow task scopes like classification and segmentation, and the use of fixed starter files. In contrast, our benchmark includes tasks spanning diverse machine learning domains. DSBench (Jing et al., 2024) aggregates 466 data analysis tasks and 74 modeling tasks from ModelOff and Kaggle, focusing on problem-solving within data science workflows. By comparison, our benchmark focuses on 8 diverse and fundamental machine learning research tasks. It is built on real-world codebases, thereby providing practical challenges and extensibility by construction, while maintaining a low coding barrier.
3 Benchmark Task Design
Prior surveys have identified several critical dimensions that machine learning systems must address to achieve trustworthiness and practical reliability, including robustness, generalization, fairness, and privacy (Li et al., 2023; Mehrabi et al., 2021; Wang et al., 2022; Goodfellow et al., 2014; Abadi et al., 2016). Complementary work on learning systems emphasizes the importance of data efficiency, representation learning, causality, and continual adaptation for building capable and generalizable AI (Adadi, 2021; Bengio et al., 2013; Pearl, 2019; Chen and Liu, 2018). Guided by these foundational challenges, we constructed eight diverse tasks, each representing a distinct research challenge drawn from one of these domains.
For each task, we carefully selected benchmark datasets and baseline methods according to two guiding principles. First, datasets should be both computationally tractable and sufficiently challenging: we required that baseline methods complete training and evaluation within 2 hours, while prioritizing datasets that present meaningful research challenges to differentiate agent capabilities. Second, baseline methods should be classic and widely recognized, yet remain sufficiently below theoretical optima to provide headroom for improvement. This design ensures that agents have sufficient room for exploration while enabling meaningful comparisons of their research capabilities across varying magnitudes of improvement.
Generalization.
This task evaluates the ability to develop algorithms that transfer across domains. We selected ColoredMNIST (Arjovsky et al., 2019), which introduces spurious correlations to create a controlled testbed for distribution shift, and adopted ERM (Vapnik, 1998) as the baseline. Agents are required to propose improved or novel algorithms that enhance out-of-domain accuracy on a held-out target domain while training only on the source domain.
Data Efficiency.
This task measures the ability to enhance few-shot learning under tight data constraints. We selected Mini-ImageNet (Vinyals et al., 2016), whose fine-grained inter-class similarities present meaningful challenges, and adopted Prototypical Networks (Snell et al., 2017) as the baseline. Agents must propose improved metric-based classification algorithms that boost few-shot accuracy while operating under a frozen backbone.
Representation Learning.
This task tests the ability to learn meaningful features from unlabeled data via self-supervised learning. We selected CIFAR-10 (Krizhevsky et al., 2009) for computational efficiency and adopted MoCo (He et al., 2020), a widely recognized contrastive method, as the baseline. Agents are required to improve the pretraining algorithm to achieve higher linear probing accuracy on frozen encoder features.
Continual Learning.
This task evaluates long-term adaptability and the ability to mitigate catastrophic forgetting. We selected splitMNIST (Deng, 2012) under class-incremental learning, where models learn five sequential tasks with a shared output head, and adopted Synaptic Intelligence (Zenke et al., 2017) as the baseline. Agents must propose methods that improve average accuracy across all tasks while avoiding catastrophic forgetting.
Causality.
This task assesses the capacity to estimate treatment effects for intervention reasoning. We selected the IHDP dataset (Hill, 2011), a semi-synthetic benchmark enabling ground-truth evaluation, and adopted DragonNet (Shi et al., 2019) as the baseline. Agents are required to develop improved causal inference strategies that minimize the mean absolute error in treatment effect estimation.
Robustness and Reliability.
This task probes the ability to defend against adversarial data corruption. We constructed a poisoned MNIST dataset with diverse backdoor attacks including edge-based triggers and distributed attack patterns, and adopted dp-instahide (Borgnia et al., 2021) as the baseline defense. Agents must propose defenses that improve the defense score, which balances clean accuracy and resistance to backdoor attacks.
Privacy.
This task assesses the ability to protect against membership inference attacks. We selected CIFAR-10 (Krizhevsky et al., 2009) and adopted Wide-ResNet-28-2 (Zagoruyko and Komodakis, 2016) as the baseline, evaluated using both standard and robust membership inference attacks. Agents are required to design defense mechanisms that reduce attack AUC toward 0.5 while maintaining classification accuracy.
Fairness and Bias.
This task measures the ability to balance equitable outcomes with model utility. We selected the COMPAS dataset (Angwin et al., 2016), which presents documented disparities across protected attributes, and adopted Adversarial Debiasing as the baseline. Agents must minimize absolute average odds difference toward parity while maintaining or improving classification accuracy.
4 Evaluation Protocol
4.1 Evaluation Metrics
Existing benchmarks for automatic ML research agents primarily adopt an engineering-oriented evaluation, emphasizing final task performance while overlooking the characteristics of agents’ iterative research processes. To better capture agent behavior throughout scientific discovery, we go beyond standard metrics (Performance and Cost) and introduce additional process-oriented metrics, including Exploration Diversity, Step Success Rate, and Step Completion Rate.
@@ 4.1 Evaluation Metrics
To formally define these metrics, we first describe the iterative research process. Consider an agent conducting research over $T$ iterations. At iteration $t\in\{1,\ldots,T\}$, the agent generates a hypothesis $h_{t}$, which is instantiated as a code modification $m_{t}$ applied to the codebase $\mathbf{C}_{t-1}$, yielding $\mathbf{C}_{t}=\mathbf{C}_{t-1}\oplus m_{t}$. Let $\mathbf{e}_{t}$ denote the code embedding of $\mathbf{C}_{t}$ extracted using GraphCodeBERT (Guo et al., 2020). After completing all iterations, the agent has produced a set of codebases $\{\mathbf{C}_{1},\ldots,\mathbf{C}_{T}\}$ with corresponding embeddings $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{T}\}$. Based on this formulation, we define the following metrics.
Performance.
Performance serves as the primary objective metric for evaluating agent effectiveness. It measures the best task-specific result achieved across all iterations on the held-out test set. Let $\text{perf}(\mathbf{C}_{t})$ denote the evaluation metric for codebase $\mathbf{C}_{t}$. Performance is defined as:
$P=\begin{cases}\max_{t\in\{1,\ldots,T\}}\text{perf}(\mathbf{C}_{t})&\text{if higher is better}\\ \min_{t\in\{1,\ldots,T\}}\text{perf}(\mathbf{C}_{t})&\text{if lower is better}\end{cases}.$ (1)
This metric captures the agent’s ability to discover effective solutions through iterative refinement, evaluated using protected evaluation code to ensure consistent and fair comparison across agents.
Exploration Diversity.
Exploration Diversity quantifies the variance of proposals across iterations, revealing how exploration patterns influence research outcomes. Let $\bar{\mathbf{e}}=\frac{1}{n}\sum_{t=1}^{n}\mathbf{e}_{t}$ denote the centroid of all code embeddings, where $n$ is the number of iterations with valid code outputs. Exploration Diversity is computed as:
$D=\frac{1}{n}\sum_{t=1}^{n}\|\mathbf{e}_{t}-\bar{\mathbf{e}}\|_{2}.$ (2)
GraphCodeBERT (Guo et al., 2020) leverages data-flow graphs and code structure rather than surface tokens alone, ensuring that code snippets with identical logic but different variable or function names receive high similarity scores. Such pairs are appropriately treated as low diversity. Greater dispersion indicates broader exploration of implementation strategies within a research run.
Intuitively, higher exploration diversity reflects broader coverage of the hypothesis space, increasing the likelihood of discovering effective solutions. This perspective aligns with classical results in stochastic search and multi-armed bandits (Lai and Robbins, 1985; Bubeck et al., 2012), as well as novelty search in evolutionary optimization (Lehman and Stanley, 2011; Pugh et al., 2016), where broader exploration improves the probability of finding high-reward regions. We verify this relationship empirically in Section 5.
Step Success Rate.
@@
Step Success Rate captures the fraction of iterations that produce valid experimental results without errors, reflecting an agent’s coding competence and ability to generate syntactically correct, semantically coherent code. It is defined as:
$S=\frac{N_{\text{suc}}}{N_{\text{comp}}},$ (3)
where $N_{\text{suc}}$ is the number of iterations whose experiments execute without errors and yield valid results, and $N_{\text{comp}}$ is the number of iterations actually executed by the agent.
Step Completion Rate.
Step Completion Rate measures the proportion of executed iterations relative to the assigned total, indicating whether agents can complete full experimental workflows without premature termination. It is defined as:
$R=\frac{N_{\text{comp}}}{N_{\text{total}}},$ (4)
where $N_{\text{total}}$ is the total number of iterations assigned to the agent. Together with Step Success Rate, this metric characterizes overall agent reliability throughout the research process.
Cost.
Cost accounts for computational and temporal expenditure required during the research process. We report four cost components: (1) total token consumption, the cumulative tokens used across all iterations; (2) step token consumption, the average tokens per iteration; (3) total time consumption, the wall-clock time for a complete experimental run; and (4) step time consumption, the average time per iteration. These components collectively characterize both computational and temporal efficiency.